# Generalized Linear Bandits: Almost Optimal Regret with One-Pass Update

### v1

@@ A.1 Main Proof
$+\frac{2\eta RSC_{\mu}}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}.$ (11)
@@
with probability at least $1-\delta$. Next, we proceed to analyze $\mathtt{term~{}(b)}$. Under the condition that $\lambda\geq\max\{14d\eta R^{2},2C_{\mu}/g(\tau)\}$, Lemma 6 with $\alpha=\frac{3}{2}\eta$ shows that
$\mathtt{term~{}(b)}\leq\frac{1}{3\eta}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+3d\eta\ln\bigg{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\bigg{)}.$ (14)
@@
$\lVert\theta_{t+1}-\theta_{*}\rVert^{2}_{H_{t}}\leq{}$ $4\lambda S^{2}+2\eta\log\Big{(}\frac{1}{\delta}\Big{)}+6d\eta^{2}\ln\bigg{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\bigg{)}$
$+\frac{2\eta RSC_{\mu}}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\frac{1}{3}\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}$
[…]
where the last inequality holds due to the condition $\lambda\geq 6\eta RSC_{\mu}/g(\tau)$. ∎
@@ A.2 Useful Lemmas
$+\frac{C_{\mu}(2+2RS)}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}.$
Furthermore, if $\lambda\geq 6\eta C_{\mu}RS/g(\tau)$, we can further have
@@
[…]
@@
$\langle\nabla\ell_{s}(\theta_{s+1})-\nabla\widetilde{\ell}_{s}(\theta_{s+1}),\theta_{s+1}-\theta_{*}\rangle={}$ $\frac{\mu^{\prime\prime}(X_{s}^{\top}\bm{\xi}_{s})}{2g(\tau)}\cdot\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{X_{s}X_{s}^{\top}}\cdot X_{s}^{\top}(\theta_{s+1}-\theta_{*})$
[…]
where $\bm{\xi}_{s}\in\Theta$ lies on the line connecting $\theta_{s}$ and $\theta_{s+1}$. The first equality follows from the definition of ~$\nabla\widetilde{\ell}_{s}(\theta_{s+1})$ and the Taylor expansion with Lagrange’s remainder. The first inequality uses the self-concordance property of the loss function, which ensures that $\mu^{\prime\prime}(z)\leq R\mu^{\prime}(z)$ and $\mu^{\prime}(z)\leq C_{\mu}$ for all $z\in[-S,S]$. The last inequality is due to $\lVert X_{s}\rVert_{2}\leq 1$ and $\lVert\theta_{s}-\theta_{s+1}\rVert_{2}\leq 2S$.
@@
[…]
@@
[…]
@@
Under Assumption 12 and 3, let $P_{s}=\mathcal{N}(\theta_{s},\alpha H_{s}^{-1})$ be a Gaussian distribution with mean $\theta_{s}$ and covariance matrix $\alpha H_{s}^{-1}$, where $H_{s}=\lambda I_{d}+\sum_{\tau=1}^{s-1}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$ and $\alpha>0$ is any positive constant. We denote by $\theta_{s}$ the model returned by (3). Then, setting $\lambda\geq\max\{64d\alpha R^{2}/7,2C_{\mu}/g(\tau)\}$ we have
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+2d\alpha\ln\bigg{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\bigg{)},$
@@
for any distribution $Q_{s}$ defined over $\mathbb{R}^{d}$, where $P_{s}^{*}(\theta)\propto P_{s}(\theta)\cdot e^{-\ell_{s}(\theta)}$ for all $\theta\in\mathbb{R}^{d}$. Here, we choose $Q_{s}=\mathcal{N}(\theta_{s+1},\alpha H_{s}^{-1})$ as a Gaussian distribution with mean $\theta_{s+1}$ and covariance $\alpha H_{s}^{-1}$.
@@
[…]
where $\mathcal{D}_{\ell_{s}}(\theta,\theta_{s+1})=\ell_{s}(\theta)-\ell_{s}(\theta_{s+1})-\langle\nabla\ell_{s}(\theta_{s+1}),\theta-\theta_{s+1}\rangle$ is the Bregman divergence of $\ell_{s}$ between $\theta$ and $\theta_{s+1}$ and $M_{s}=\frac{\lambda}{2}I_{d}+\sum_{\tau=1}^{s}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$. In the above, the second equality follows from the definition of $Q_{s}$. The final inequality follows from Lemma 7 in Appendix C, since $\ell_{s}$ is self-concordant and the condition $\lambda\geq\max\{64d\alpha R^{2}/7,2C_{\mu}/g(\tau)\}$ holds.
@@
$\mathtt{term~{}(b)}=\frac{1}{2\alpha}\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{H_{s}}.$ (23)
@@
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+2\alpha\ln\left(\frac{\mbox{det}(M_{t})}{\mbox{det}(\lambda I_{d}/2)}\right).$
The determinant of the matrix $M_{t}$ can be further bounded by
$\det(M_{t})=\det\bigg{(}\lambda I_{d}+\sum_{s=1}^{t-1}\nabla^{2}\ell_{s}(\theta_{s+1})\bigg{)}\leq\mbox{det}\Big{(}\big{(}\lambda+C_{\mu}t/g(\tau)\big{)}I_{d}\Big{)}\leq\big{(}\lambda+C_{\mu}t/g(\tau)\big{)}^{d}.$
@@
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+2d\alpha\ln\Big{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\Big{)},$
@@
Let $\ell_{s}(\theta)$ be the loss function of the maximum likelihood estimator and $Q_{s}=\mathcal{N}(\theta_{s+1},\alpha H_{s}^{-1})$ be a Gaussian distribution with mean $\theta_{s+1}$ and covariance matrix $\alpha H^{-1}_{s+1}$ with $H_{s}=\lambda I_{d}+\sum_{\tau=1}^{s-1}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$. Under Assumption 12 and 3 and setting $\lambda\geq\max\{64d\alpha R^{2}/7,2C_{\mu}/g(\tau)\}$, for any constant $\alpha>0$, we have
$\mathbb{E}_{\theta\sim Q_{s}}[\mathcal{D}_{\ell_{s}}(\theta,\theta_{s+1})]\leq 2\alpha\Big{(}\log\mbox{det}(M_{s})-\log\mbox{det}(M_{s-1})\Big{)},$
where $M_{s}=\frac{\lambda}{2}I_{d}+\sum_{\tau=1}^{s}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$.
@@
where $\widetilde{h}_{s}(\theta)=\int_{v=0}^{1}(1-v)\mu^{\prime}\big{(}\theta_{s+1}+v(\theta-\theta_{s+1})\big{)}\,\mathrm{d}v\cdot X_{s}X_{s}^{\top}/g(\tau)$. According to Lemma 8 in Appendix C and the condition $\lVert X_{t}\rVert_{2}\leq 1$ by Assumption 1, we can bound the Hessian $\widetilde{h}_{s}(\theta)$ by
@@
[…]
@@
where $\lambda_{i}$ is the $i$-th largest eigenvalue of the matrix $\big{(}\nabla^{2}\ell_{s}(\theta_{s+1})\big{)}^{\frac{1}{2}}H_{s}^{-1}\big{(}\nabla^{2}\ell_{s}(\theta_{s+1})\big{)}^{\frac{1}{2}}$ and $\mathbf{e}_{i}$ is a set of orthogonal basis. Then, we have
@@
[…]
where $\mbox{Tr}(A)$ denotes the trace of matrix $A$. In the above, the last inequality is due to the fact that $\mathbb{E}_{X_{i},X_{j}\sim\mathcal{N}(0,1)}[X_{i}X_{j}]=3$. The last inequality is due to $\mbox{trace}(AB)=\mbox{trace}(BA)$ for matrix $A,B\in\mathbb{R}^{d\times d}$. Recall that $H_{s}=\lambda I_{d}+\sum_{\tau=1}^{s-1}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$. Define $M_{s}=\frac{\lambda}{2}I_{d}+\sum_{\tau=1}^{s}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$. Given the constraint $\lambda\geq 2C_{\mu}/g(\tau)$, one can verify that $H_{s}\succcurlyeq M_{s}$. Then we have
$\mbox{Tr}\Big{(}H_{s}^{-1}\big{(}\nabla^{2}\ell_{s}(\theta_{s+1})\big{)}\Big{)}\leq{}$ $\mbox{Tr}\Big{(}M_{s}^{-1}\big{(}\nabla^{2}\ell_{s}(\theta_{s+1})\big{)}\Big{)}=\mbox{Tr}\Big{(}M_{s}^{-1}\big{(}M_{s}-M_{s-1})\Big{)}$
[…]

### v2

@@ A.1 Main Proof
$+\frac{2\eta RSL_{\mu}}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}.$ (11)
@@
with probability at least $1-\delta$. Next, we proceed to analyze $\mathtt{term~(b)}$. Under the condition that $\lambda\geq 14d\eta R^{2}$, Lemma 6 with $\alpha=\frac{3}{2}\eta$ shows that
$\mathtt{term~(b)}\leq\frac{1}{3\eta}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+d(3\eta+\frac{1}{2})\ln\bigg(1+\frac{L_{\mu}t}{\lambda g(\tau)}\bigg).$ (14)
@@
$\lVert\theta_{t+1}-\theta_{*}\rVert^{2}_{H_{t}}\leq{}$ $4\lambda S^{2}+2\eta\log\Big(\frac{1}{\delta}\Big)+d\eta(6\eta+1)\ln\bigg(1+\frac{L_{\mu}t}{\lambda g(\tau)}\bigg)$
$+\frac{2\eta RSL_{\mu}}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\frac{1}{3}\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}$
$\leq{}$ $4\lambda S^{2}+2\eta\log\Big(\frac{1}{\delta}\Big)+d\eta(6\eta+1)\ln\bigg(1+\frac{L_{\mu}t}{\lambda g(\tau)}\bigg),$
where the last inequality holds due to the condition $\lambda\geq 6\eta RSL_{\mu}/g(\tau)$. ∎
@@ A.2 Useful Lemmas
$+\frac{L_{\mu}(2+2RS)}{g(\tau)}\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}-\sum_{s=1}^{t}\|\theta_{s}-\theta_{s+1}\|^{2}_{H_{s}}.$
Furthermore, if $\lambda\geq 6\eta L_{\mu}RS/g(\tau)$, we can further have
@@
$={}$ $\langle\nabla\ell_{s}(\theta_{s+1})-\nabla\widetilde{\ell}_{s}(\theta_{s+1}),\theta_{s+1}-\theta_{*}\rangle+\langle\nabla\widetilde{\ell}_{s}(\theta_{s+1}),\theta_{s+1}-\theta_{*}\rangle$
$\leq{}$ $\langle\nabla\ell_{s}(\theta_{s+1})-\nabla\widetilde{\ell}_{s}(\theta_{s+1}),\theta_{s+1}-\theta_{*}\rangle$
@@
$\langle\nabla\ell_{s}(\theta_{s+1})-\nabla\widetilde{\ell}_{s}(\theta_{s+1}),\theta_{s+1}-\theta_{*}\rangle$
$={}$ $\langle\nabla\ell_{s}(\theta_{s+1})-\nabla\ell_{s}(\theta_{s})-\nabla^{2}\ell_{s}(\theta_{s})(\theta_{s+1}-\theta_{s}),\theta_{s+1}-\theta_{*}\rangle$
$={}$ $\frac{1}{g(\tau)}\langle\mu(X_{s}^{\top}\theta_{s+1})\cdot X_{s}-\mu(X_{s}^{\top}\theta_{s})\cdot X_{s}-\mu^{\prime}(X_{s}^{\top}\theta_{s})\cdot X_{s}X_{s}^{\top}(\theta_{s+1}-\theta_{s}),\theta_{s+1}-\theta_{*}\rangle$
$={}$ $\frac{\mu^{\prime\prime}(X_{s}^{\top}\bm{\xi}_{s})}{2g(\tau)}\cdot\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{X_{s}X_{s}^{\top}}\cdot X_{s}^{\top}(\theta_{s+1}-\theta_{*})$
$\leq{}$ $\frac{R}{2g(\tau)}\cdot\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{\mu^{\prime}(X_{s}^{\top}\bm{\xi}_{s})X_{s}X_{s}^{\top}}\cdot X_{s}^{\top}(\theta_{s+1}-\theta_{*})$
$\leq{}$ $RS\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{\nabla^{2}\ell_{s}(\bm{\xi}_{s})}$
$\leq{}$ $\frac{RSL_{\mu}}{g(\tau)}\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{2},$ (19)
where $\bm{\xi}_{s}\in\Theta$ lies on the line connecting $\theta_{s}$ and $\theta_{s+1}$. The first equality follows from the definition of $\nabla\widetilde{\ell}_{s}(\theta_{s+1})$ and the Taylor expansion with Lagrange’s remainder. The first inequality uses the self-concordance property of the loss function, which ensures that $\mu^{\prime\prime}(z)\leq R\mu^{\prime}(z)$ and $\mu^{\prime}(z)\leq L_{\mu}$ for all $z\in[-S,S]$. The last inequality is due to $\lVert X_{s}\rVert_{2}\leq 1$ and $\lVert\theta_{s}-\theta_{s+1}\rVert_{2}\leq 2S$.
@@
$\leq{}$ $\frac{1}{2+2RS}\left(\lVert\theta_{1}-\theta_{*}\rVert^{2}_{H_{1}}-\lVert\theta_{t+1}-\theta_{*}\rVert^{2}_{H_{t+1}}-\sum_{s=1}^{t}\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{H_{s}}\right)+\frac{RSL_{\mu}}{g(\tau)}\lVert\theta_{s}-\theta_{s+1}\rVert_{2}^{2}.$
@@
$={}$ $M_{t-1}\cdot\int\frac{\mathbb{E}_{\theta\sim P_{t}}[p(r|X_{t}^{\top}\theta)]}{p(r|X_{t}^{\top}\theta_{*})}\cdot p(r|X_{t}^{\top}\theta_{*})\mathrm{d}r$
$={}$ $M_{t-1}\cdot\int\mathbb{E}_{\theta\sim P_{t}}[p(r|X_{t}^{\top}\theta)]\mathrm{d}r$
$={}$ $M_{t-1}\cdot\mathbb{E}_{\theta\sim P_{t}}\left[\int p(r|X_{t}^{\top}\theta)\mathrm{d}r\right]$
$={}$ $M_{t-1},$
@@
Under Assumption 12 and 3, let $P_{s}=\mathcal{N}(\theta_{s},\alpha H_{s}^{-1})$ be a Gaussian distribution with mean $\theta_{s}$ and covariance matrix $\alpha H_{s}^{-1}$, where $H_{s}=\lambda I_{d}+\sum_{\tau=1}^{s-1}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$ and $\alpha$ is any positive constant. We denote by $\theta_{s}$ the model returned by (3). Then, setting $\lambda\geq 64d\alpha R^{2}/7$ we have
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+d(2\alpha+\frac{1}{2})\ln\bigg(1+\frac{2L_{\mu}t}{\lambda g(\tau)}\bigg),$
@@
for any distribution $Q_{s}$ defined over $\mathbb{R}^{d}$, where $P_{s}^{*}(\theta)\propto P_{s}(\theta)\cdot e^{-\ell_{s}(\theta)}$ for all $\theta\in\mathbb{R}^{d}$. Here, we choose $Q_{s}=\mathcal{N}(\theta_{s+1},\alpha H_{s+1}^{-1})$ as a Gaussian distribution with mean $\theta_{s+1}$ and covariance $\alpha H_{s+1}^{-1}$.
@@
$={}$ $\ell_{s}(\theta_{s+1})+\mathbb{E}_{\theta\sim Q_{s}}[\mathcal{D}_{\ell_{s}}(\theta,\theta_{s+1})]$
$\leq{}$ $\ell_{s}(\theta_{s+1})+2\alpha\Big(\log\mbox{det}(H_{s+1})-\log\mbox{det}(H_{s})\Big),$ (22)
where $\mathcal{D}_{\ell_{s}}(\theta,\theta_{s+1})=\ell_{s}(\theta)-\ell_{s}(\theta_{s+1})-\langle\nabla\ell_{s}(\theta_{s+1}),\theta-\theta_{s+1}\rangle$ is the Bregman divergence of $\ell_{s}$ between $\theta$ and $\theta_{s+1}$. In the above, the second equality follows from the definition of $Q_{s}$. The last line follows from Lemma 7 in Appendix C, since $\ell_{s}$ is self-concordant and the condition $\lambda\geq 64d\alpha R^{2}/7$ holds.
@@
$\mathtt{term~(b)}={}$ $\frac{1}{2}\left(\log\mbox{det}(H_{s+1})-\log\mbox{det}(H_{s})+\mbox{Tr}\left(H_{s}H_{s+1}^{-1}\right)+\frac{\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{H_{s}}}{\alpha}-d\right)$
$\leq{}$ $\frac{1}{2}\left(\log\mbox{det}(H_{s+1})-\log\mbox{det}(H_{s})\right)+\frac{1}{2\alpha}\lVert\theta_{s}-\theta_{s+1}\rVert^{2}_{H_{s}}.$ (23)
@@
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+(2\alpha+\frac{1}{2})\ln\left(\frac{\mbox{det}(H_{t+1})}{\mbox{det}(\lambda I_{d})}\right).$
The determinant of the matrix $H_{t+1}$ can be further bounded by
$\det(H_{t+1})=\det\bigg(\lambda I_{d}+\sum_{s=1}^{t}\nabla^{2}\ell_{s}(\theta_{s+1})\bigg)\leq\mbox{det}\Big(\big(\lambda+L_{\mu}t/g(\tau)\big)I_{d}\Big)\leq\Bigl(\lambda+L_{\mu}t/g(\tau)\Bigr)^{d}.$
@@
$\sum_{s=1}^{t}m_{s}(P_{s})\leq{}$ $\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})+\frac{1}{2\alpha}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+d(2\alpha+\frac{1}{2})\ln\Big(1+\frac{L_{\mu}t}{\lambda g(\tau)}\Big),$
@@
Let $\ell_{s}(\theta)$ be the loss function of the maximum likelihood estimator and $Q_{s}=\mathcal{N}(\theta_{s+1},\alpha H_{s+1}^{-1})$ be a Gaussian distribution with mean $\theta_{s+1}$ and covariance matrix $\alpha H_{s+1}^{-1}$ with $H_{s}=\lambda I_{d}+\sum_{\tau=1}^{s-1}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$. Under Assumption 12 and 3 and setting $\lambda\geq 64d\alpha R^{2}/7$, for any constant $\alpha>0$, we have
$\mathbb{E}_{\theta\sim Q_{s}}[\mathcal{D}_{\ell_{s}}(\theta,\theta_{s+1})]\leq 2\alpha\Big(\log\mbox{det}(H_{s+1})-\log\mbox{det}(H_{s})\Big),$
@@
where $\widetilde{h}_{s}(\theta)=\int_{v=0}^{1}(1-v)\mu^{\prime}\big(X_{s}^{\top}\theta_{s+1}+vX_{s}^{\top}(\theta-\theta_{s+1})\big)\,\mathrm{d}v\cdot X_{s}X_{s}^{\top}/g(\tau)$. According to Lemma 8 in Appendix C and the condition $\lVert X_{t}\rVert_{2}\leq 1$ by Assumption 1, we can bound the Hessian $\widetilde{h}_{s}(\theta)$ by
@@
$\leq{}$ $\sqrt{\mathbb{E}_{\theta\sim Q_{s}}\bigg[{e^{2R^{2}\lVert\theta-\theta_{s+1}\rVert_{2}^{2}}}\bigg]\cdot\mathbb{E}_{\theta\sim Q_{s}}\bigg[\lVert\theta-\theta_{s+1}\rVert^{4}_{\nabla^{2}\ell_{s}(\theta_{s+1})}\bigg]}$
$\leq{}$ $\sqrt{\frac{4}{3}\mathbb{E}_{\theta\sim Q_{s}}\bigg[\Big\|\big(\nabla^{2}\ell_{s}(\theta_{s+1})\big)^{\frac{1}{2}}(\theta-\theta_{s+1})\Big\|^{4}_{2}\bigg]},$ (25)
@@
where $\lambda_{i}$ is the $i$-th largest eigenvalue of the matrix $\big(\nabla^{2}\ell_{s}(\theta_{s+1})\big)^{\frac{1}{2}}H_{s+1}^{-1}\big(\nabla^{2}\ell_{s}(\theta_{s+1})\big)^{\frac{1}{2}}$ and $\mathbf{e}_{i}$ is a set of orthogonal basis. Then, we have
@@
$={}$ $\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{d}\alpha^{2}\lambda_{i}\lambda_{j}\mathbb{E}_{X_{i},X_{j}\sim\mathcal{N}(0,1)}[X_{i}^{2}X_{j}^{2}]}$
$={}$ $\sqrt{3}\alpha\mbox{Tr}\Big(H_{s+1}^{-1}\big(\nabla^{2}\ell_{s}(\theta_{s+1})\big)\Big).$ (26)
where $\mbox{Tr}(A)$ denotes the trace of matrix $A$. In the above, the last inequality is due to the fact that $\mathbb{E}_{X_{i},X_{j}\sim\mathcal{N}(0,1)}[X_{i}X_{j}]=3$. The last inequality is due to $\mbox{trace}(AB)=\mbox{trace}(BA)$ for matrix $A,B\in\mathbb{R}^{d\times d}$. Recall that $H_{s+1}=\lambda I_{d}+\sum_{\tau=1}^{s}\nabla^{2}\ell_{\tau}(\theta_{\tau+1})$.
$\mbox{Tr}\Big(H_{s+1}^{-1}\big(\nabla^{2}\ell_{s}(\theta_{s+1})\big)\Big)\leq{}$ $\mbox{Tr}\Big(H_{s+1}^{-1}(H_{s+1}-H_{s})\Big)=\mbox{Tr}\Big(I-H_{s}H_{s+1}^{-1}\Big)$
$\leq{}$ $\log\mbox{det}(H_{s+1})-\log\mbox{det}(H_{s}).$ (27)
@@
A.3 Computational Cost Discussion
This part discusses the time and space complexity of solving the optimization problem (3).
Proposition 1.
Proof of Proposition 1.
We begin with the analysis on the time complexity, followed by the discussion on the space complexity.
Space Complexity. Regarding space complexity, it suffices to store the current model $\theta_{t}$, the gradient $\nabla\ell_{t}(\theta_{t})$, the inverse Hessian matrix $H_{t}^{-1}$, and $\widetilde{H}_{t}^{-1}$ throughout the optimization process, resulting in a total space complexity of $\mathcal{O}(d^{2})$. ∎